OpenStudy (anonymous):

If sin theta=3/5 and theta is in quad 2 the exact form of cos (5pi/3-theta) is .....? I just need help with it I can do the work but a little help could help lol

OpenStudy (anonymous):

o.O thats not the same question

OpenStudy (anonymous):

Yeah I know for sure thats not the right answer lol

OpenStudy (anonymous):

Use the compound angle formula for cos,

OpenStudy (0_0youscareme):

i put the link there bcuz that he just put copy and paste

OpenStudy (anonymous):

this thing is refusing to let me use the equation tool... cos(A-B)=cosAcosB+sinAsinB

OpenStudy (0_0youscareme):

i never said it had the answer

OpenStudy (anonymous):

Nono lol I was talking about jacob

OpenStudy (0_0youscareme):

yeah but i put the link were he copy and pasted the answer some1 else put

OpenStudy (anonymous):

I got the answer 3sqrt3-4/10 but thats the same answer as another problem I did

OpenStudy (anonymous):

he deleted it i think... XD

OpenStudy (anonymous):

\[\sin \Theta=\frac{ 5 }{ 3 }\] \[\cos \theta=\frac{ 4 }{ 5 }\] \[\cos (\frac{ 5\pi }{ 3 }-\theta)=\cos \frac{ 5\pi }{ 3 } \cos \theta + \sin\]

OpenStudy (anonymous):

oops incomplete... k im not used to this equation tool. meant to be sin5pi/3sin(theta)

OpenStudy (anonymous):

@zepdrix Do you have any idea of how this is done? I think my work is correct

OpenStudy (anonymous):

\[=\frac{ 1 }{ 2 }\frac{ 4 }{ 5 }+(-\frac{ \sqrt{3} }{ 4 })\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

I think I got it now after doing it

OpenStudy (anonymous):

But the way you were doing it was a little wrong lol